OpenStudy (anonymous):

Can someone explain the steps to get the answer of the function f(x+1) =

OpenStudy (anonymous):

sub x+1 into the variable of the function rule and simplify.

OpenStudy (anonymous):

do you have a specific function

OpenStudy (anonymous):

yes , function: f(x) =3x+2

OpenStudy (anonymous):

replace x+1 in the variabl ein the function f f(x+1)=3(x+1)+2 now simplify f(x+1)=3x+3+2 or f(x+1)=3x+5

OpenStudy (anonymous):

i got : 3(x+1)+2 3x +5 then i couldnt solve anymore and i got stuck

OpenStudy (anonymous):

hello